# Quantifying Self-Awareness of Knowledge in Large Language Models

### v1

@@ Abstract
Hallucination prediction in large language models (LLMs) is often interpreted as a sign of self-awareness. However, we argue that such performance can arise from question-side shortcuts rather than true model-side introspection. To disentangle these factors, we propose the Approximate Question-side Effect (AQE), which quantifies the contribution of question-awareness. Our analysis across multiple datasets reveals that much of the reported success stems from exploiting superficial patterns in questions. We further introduce SCAO (Semantic Compression by Answering in One word), a method that enhances the use of model-side signals. Experiments show that SCAO achieves strong and consistent performance, particularly in settings with reduced question-side cues, highlighting its effectiveness in fostering genuine self-awareness in LLMs. The code is available online (https://github.com/ybseo-ac/AQE).
@@ 1 Introduction
The defining and quantifying of the human-like mental attributes of large language models (LLMs) lie at the heart of a long-standing human curiosity: whether artificial systems can possess minds akin to our own. While the field has demonstrated that LLMs can rival or even surpass human-level performance in rational reasoning (e.g., reading, solving a logical problem) [4, 11, 29], higher-order human-like traits (e.g., self-awareness, personality, emotion) remain poorly understood, despite growing interest [21, 41], partly due to the inherent ambiguity in their definition and measurement.
Among higher-order attributes, self-awareness of knowledge has been considered as a key topic, given its strong association with hallucination detection—a critical factor in ensuring the reliability of LLMs. While hallucination arises from various causes (e.g., failure of reading, reasoning), it is well known that a major cause is when the model is asked questions on the knowledge not pre-trained [37]. Humans possess an innate ability of self-awareness to recognize the possession of their knowledge [13, 17], and thus refuse to answer when they lack the relevant knowledge. In contrast, LLMs lack such self-awareness and therefore tend to produce any probable output even when they lack knowledge, which leads to hallucination. In this respect, self-awareness is a fundamental component in achieving reliable language generation.
@@
Then, how can we define and measure self-awareness of LLMs? Prior work has often equated self-awareness with the behaviour of hallucination detection itself, perhaps because it is the ultimate practical goal we expect from self-awareness. Indeed, recent studies have reported surprisingly high precision on hallucination detection [35, 42, 26, 13, 1], raising hopes that self-awareness in LLMs may be within reach.
@@
However, in this work, we argue that hallucination prediction (i.e., early detection) does not equate to self-awareness. This is because two distinct sources of information are typically involved in the prediction process: (1) information about the model itself and (2) information about the question. As such, hallucination prediction reflects a mixture of model-awareness (self-awareness) and question-awareness. To isolate the true contribution of self-awareness, we propose to disentangle these two components. Through this separation, we offer a rigorous and measurable definition of self-awareness, which we believe serves as a foundational step toward investigating the higher-order mental faculties of LLMs. To enable this separation, we introduce a metric to quantify the question-awareness in hallucination detection, the Approximate Question-side Effect (AQE). The contribution of self-awareness is estimated by subtracting AQE from hallucination detection precision.
Quantifying the contribution of self-awareness to hallucination prediction performance is not merely a matter of curiosity: it carries significant practical implications, in our observation in §3.2. When hallucination prediction relies on question-awareness, models tend to exploit dataset-specific shortcuts to easily achieve high performance. However, such performance fails to generalize when the distribution of questions shifts even slightly. In contrast, relying on self-awareness tends to produce more consistent results across varying question distributions, as it utilizes model-side information rather than question-side cues. To support this claim, in §B, we first analyze dataset-specific shortcut patterns in existing datasets used in previous hallucination detection works. Then we provide the following empirical findings in §3.3 and §5: (1) Datasets containing shortcut cues tend to exhibit higher AQE scores, suggesting stronger question-awareness signals. However, AQE scores drop significantly under shortcut-controlled settings (e.g., out-of-domain scenarios). (2) In such high-AQE datasets, hallucination detection performance shows limited generalizability
@@
Lastly, we also propose a method to enhance the use of model-side information, by leveraging the confidence scores of LLMs more effectively. The proposed method is Semantic Compression by Answering in One word (SCAO). We demonstrate that SCAO performs particularly well in low AQE settings, and the performance is consistent across varying settings of questions.
@@
  • We propose a measurable definition of “self-awareness” in LLMs, distinct from question-awareness, and introduce AQE to quantify its contribution. This offers a concrete example for studying human-like mental attributes in language models.
  • We analyze existing hallucination prediction datasets and reveal that much of the reported performance stems from question-side shortcuts, limiting generalizability.
  • We introduce SCAO, a simple one-word answering method that enhances model-side signal utilization, showing strong performance particularly when question-side shortcuts are unavailable and self-awareness becomes essential.
@@ 2 Definition
In this section, we first examine the definition of self-awareness of human. Next, we define the task formulation of hallucination prediction, to establish a definition of self-awareness in LLMs. And we review the definitions from previous works. More detailed review on the related works of neuropsychology and LLMs is in §A.
@@
Definition of self-awareness of human In psychology, human self-awareness is defined as the capability of perception of one’s own mental processes or states, which includes thoughts, feelings, emotions, and knowing [28]. In this work, we focus on the self-awareness of knowing certain pieces of knowledge, which is also referred to as self-knowledge [41].
@@
Some studies on LLMs also borrow the term “meta-cognition” from psychology to refer to self-knowledge [20], which may be inaccurate. It is because meta-cognition focuses solely on the conscious level [17], while recent cognitive neuroscience research suggests that the human brain conducts judgment of knowing even at the unconscious level [13], which is even before consciously recognizing the meaning of the question.
@@
Definition of hallucination prediction The term “hallucination” has been broadly used to refer to the phenomenon where a model provides an incorrect answer to a given question [18, 26, 5]. Thus, “hallucination detection” refers to the task of predicting whether a response is incorrect [18, 6].
@@
In this work, we focus specifically on (1) hallucination from factoid questions that examine whether the model possesses certain knowledge, as this has been widely used as a clear and straightforward scenario for exploring hallucination detection [35, 42]. (2) And we focus on an early detection (i.e., prediction) scenario [35, 6, 1] where $k$ is predicted before answer generation, as this setting is more appropriate for examine self-awareness (we describe this in the next section).
@@
To formulate the common problem setting of hallucination prediction, let $\theta$ represent the model, $x$ the query, and $y$ the answer label. The model $\theta$ infers $\hat{y}$ from the input $x$. The correctness of $\hat{y}$ can then be measured by evaluating its similarity to $y$, denoted as $k$, representing a binary value (True/False). Common evaluation methods include string matching [42], GLUE, and G-eval [24]. During this process, a datapoint $s_{\theta,x}=\theta(x)$ is extracted from $\theta$, which contains information about how $\theta$ perceived $x$. We denote this as $s$ for simplicity. Through a series of question-answering and evaluation processes, we obtain a dataset $\mathcal{D}=\{(s_{i},k_{i})\}_{i=1}^{N}$, where $N$ is the dataset size. From this, hallucination prediction is defined as a binary classification task where a learnable module $\phi$ learn to take input $s_{i}$ to predict $k_{i}$, which we note $\hat{k_{i}}=\phi(s_{i})$.
@@
Assuming $\theta$ as a transformer model [38], $s$ can be mainly two forms (1) Hidden state vectors: Transformer models are composed of multiple attention layers, where each layer passes a fixed-size vector (i.e., hidden states) to the next. These vectors encode the semantic and contextual information of the input $x$ [31], and are also known to contain information about the model’s response that will be generated [19], providing a cue for hallucination prediction. (2) Confidence score: This refers to the softmax probability that the causal LLMs predict for the next token generation. While the hidden state is a high-dimensional representation (4096 dimensions in LLaMA-3-8B), the confidence score is a scalar value. As the hidden state contains richer information, it has achieved the best performance as a source for hallucination prediction and has been regarded as the main source [35, 6].
@@ 2.1 Formulating the self-awareness of LLM
We defined hallucination prediction as the process where $\phi$ perceives $s$ to infer $\hat{k}$ to predict $k$. As $s$ represents the state of the model after it has seen a question, it inherently contains information of two distinct objects, the question-side and the model-side, as illustrated in Figure 1. The question-side refers to the objective information that can be shared across different models, such as the domain of question (e.g., science, math, society) and the type of question (e.g., multiple choice, open-ended). For human, this type of information is derived from rational abilities (e.g., classification and reading comprehension), rather than from higher-order mental capacity such as self-awareness [28]. In contrast, the model-side information refers to model-specific attributes, such as the possession of the knowledge that is needed for response, or the degree of confidence for answering. In humans, this corresponds to the domain of self-awareness.
Let us denote the representation of the question-side information in $s$ as $s_{Q}$. If we exclude $s_{Q}$ from $s$, what remains is the information specific to the model $\theta$, denoted as $s_{M}$. Simplifying this, we can rewrite this as $s\approx s_{Q}\oplus s_{M}$, if $s$ is $d$-dim hidden state. By definition, we assume that $s_{Q}$ is the question-side information that is independent of model-side information ($s_{M}$), and thus the two share no overlap. For example, the domain and type of a question remain invariant regardless of the state of the model. Therefore, the previously defined hallucination prediction $\hat{k}=\phi(s)$ can be rewritten as: $\hat{k}=\phi(s_{Q},s_{M})$.
We do not claim that the hidden state $s$ is perfectly linearly separable. However, we adopt this assumption based on the following intuition: Prior studies have empirically shown that the hidden states of transformer models encode multiple properties in a linearly separable manner. For instance, Li et al. [19] demonstrates that the hidden state contains a direction associated with the attribute of “truthfulness”, therefore linearly adding this to the hidden states results in more truthful responses.
@@ 2.2 Definition from previous works
Utilizing $s$ as self-awareness Previous works implicitly regard self-awareness as the act of hallucination detection itself. In other words, the focus has been on the act of predicting $k$ from $s$, with no consideration given to the decomposition of $s$ into $s_{Q}$ and $s_{M}$. As a result, some of the hallucination detection performance reported in those works is partially overestimated by the effect of question-side shortcuts [42, 1]. We analyze such cases in §3.1.
Utilizing $s_{Q}$ as self-awareness In another case, some works define utilization of $s_{Q}$ as self-awareness, which is the opposite concept from our view. [41] defines the term “self-knowledge” as a self-awareness on possession of certain knowledge. And from this definition, they construct the dataset SelfAware to measure the self-awareness capacity. While the definition in the paper is consistent with ours, the construction of the dataset stands on the opposite definition.
@@ 3.1 Case study on question-side shortcuts
We investigate sources of question-side shortcuts in datasets used in hallucination prediction studies. We focus on short-form closed-book factoid question answering scenarios, which include ParaRel [8], Mintaka [34], HaluEval [18], HotpotQA [39], and SimpleQuestion [40]. We describe three main sources of question-side shortcuts, and further in §B.
(1) Broken question The most commonly observed problem across all datasets is insufficient annotations of question and answer pairs. This occurs when questions and answers follow a one-to-many relationship, but the annotations fail to cover them. For example, the ParaRel consists of question-answer pairs such as: “Q: What field does Daniel Bernoulli work in? A: physics”. Although Daniel Bernoulli also worked in the fields of mathematics and medicine (one-to-many relation), only one label is provided, failing to cover all possible correct answers. If an LLM responds with a correct answer but different from the given label (e.g., mathematics), it would still be classified as hallucinated (i.e., $k$ is annotated $False$), which is incorrect. If $\phi$ is trained to predict $k$ using such $\{s,k\}$ pairs, it will likely become biased toward predicting $\hat{k}=False$ whenever the domain of question is “field of work”. This means $\phi$ learns a domain classification task, not self-awareness. While this may improve prediction performance within the dataset, it lacks generalizability. It is because its performance is likely to drop only if the quality of the dataset improves, or if questions are given from unseen domains. As a study of Zhang et al. [42] reports hallucination performance on this dataset, this shortcut might have influenced the reported scores. This issue is also found in other datasets such as SimpleQuestions, which includes instances like "Q: What is a TV action show? A: Genji Tsushin Agedama", that also fail to cover various possible answers.
@@
(2) Domain Classifying the domain of a question (e.g., science, society) alone can provide a rough guess of $k$. For example, suppose a model that is extensively trained on science data but is unfamiliar with society or history domains. Since the model is likely to make more hallucination when given questions in the history domain. In this way, $k$ can be biased toward $True$ or $False$ by domains, as shown in Figure 2. In such cases, the task of predicting $k$ with $\phi$ is more a domain classification, which is question-aware and not model-aware.
(a) HotpotQA
Type $p(k=True)$
Bridge 0.6828
Comparison 0.8477
(b) Mintaka
Type $p(k=True)$
Entity 0.5508
numerical 0.4011
date 0.5968
string 0.6315
boolean 0.7283
(3) Question type Question type (e.g., short-answer, multiple-choice) also provides a strong hint for predicting $k$. The average probability of $k=True$ (denoted as $p(k=True)$) for binary-choice questions is at least 0.5 for random choice, significantly higher than for open-ended questions with 0 for random choice, as described in 1(b). In such a case, $\phi$ may learn a shortcut where it automatically predicts $k=True$ whenever it detects a binary-choice question. HotpotQA, HaluEval, and Mintaka contain such binary-choice questions.

### v2

# Quantifying Genuine Awareness in Hallucination Prediction Beyond Question-Side Shortcuts

@@ Abstract
Many works have proposed methodologies for language model (LM) hallucination detection and reported seemingly strong performance. However, we argue that the reported performance to date reflects not only a model’s genuine awareness of its internal information, but also awareness derived purely from question-side information (e.g., benchmark hacking). While benchmark hacking can be effective for boosting hallucination detection score on existing benchmarks, it does not generalize to out-of-domain settings and practical usage. Nevertheless, disentangling how much of a model’s hallucination detection performance arises from question-side awareness is non-trivial. To address this, we propose a methodology for measuring this effect without requiring human labor, Approximate Question-side Effect (AQE). Our analysis using AQE reveals that existing hallucination detection methods rely heavily on benchmark hacking. The code is available online (https://github.com/ybseo-ac/AQE).
Quantifying Genuine Awareness in Hallucination Prediction
Beyond Question-Side Shortcuts
Yeongbin Seo                   Dongha Lee                    Jinyoung Yeo thanks: Corresponding author Department of Artificial Intelligence Yonsei University {suhcrates,donalee,jinyeo}@yonsei.ac.kr
@@ 1 Introduction
The defining and quantifying of human-like mental attributes in large language models (LLMs) lie at the heart of a long-standing question: whether artificial systems can possess minds akin to our own. While recent advances show that LLMs can rival or even surpass humans in rational reasoning tasks (Brown et al., 2020; Grattafiori et al., 2024; OpenAI, 2024), higher-order traits such as self-awareness and emotion remain poorly understood, partly due to ambiguities in their definition and measurement (Li et al., 2024b; Yin et al., 2023).
Among these traits, self-awareness of knowledge is particularly important because of its close connection to hallucination detection, which is critical for the reliability of LLMs. Although hallucination can arise from various sources, a major cause is answering questions beyond the model’s pre-trained knowledge (Tonmoy et al., 2024). Humans can recognize when they lack relevant knowledge and refrain from answering (Irak et al., 2019; Koriat, 1993), whereas LLMs lack such awareness and tend to generate plausible outputs regardless, leading to hallucination.
@@
Then, how can we define and measure self-awareness of LLMs? Prior work has often equated self-awareness with hallucination detection itself, motivated by its practical importance. Indeed, recent studies report high hallucination detection performance (Snyder et al., 2024; Zhang et al., 2024; Manakul et al., 2023; Azaria and Mitchell, 2023).
@@
However, we argue that hallucination prediction does not directly measure self-awareness, because two distinct sources of information are typically involved in the prediction process: (1) information about the model itself and (2) information about the question. As such, hallucination prediction reflects a mixture of model-awareness (self-awareness) and question-awareness. To isolate self-awareness, we disentangle these two components and introduce a Shapley-based metric, the Approximate Question-side Effect (AQE), to quantify question-awareness. The contribution of self-awareness is then estimated by subtracting AQE from hallucination detection precision.
Quantifying self-awareness has important practical implications. As shown in §4.2, hallucination predictors that rely heavily on question-awareness often exploit dataset-specific shortcuts and fail to generalize under distribution shifts. In contrast, approaches grounded in model-side information yield more robust performance. We empirically support this claim through dataset analyses and experiments in §B, §4.3, and §5.
@@
Lastly, we also propose a method to enhance the use of model-side information, by leveraging the confidence scores of LLMs more effectively. The proposed method is Semantic Compression by Answering in One word (SCAO). We demonstrate that SCAO performs particularly well in low AQE settings. Though this method shows limitations in certain settings, such as long-form question answering, it provides clues to overcoming the limitations of previous probing-based methods.
@@
  • Conceptual: We disentangle hallucination detection into self-awareness and question-awareness and provide a measurable definition of self-awareness in LLMs.
  • Methodological: We introduce AQE, a Shapley-based metric to quantify question-side effects.
  • Empirical: We show that shortcut-driven, question-aware methods fail to generalize, while model-side approaches are more robust.
2 Related works
Self-awareness in humans: Insights from cognitive neuropsychology
Extensive research in cognitive psychology and neuroscience has shown that human self-awareness—particularly in the context of knowing whether one knows something—relies on layered cognitive processes. According to studies such as Koriat (1993); Irak et al. (2019); Brown et al. (2017), two key mechanisms underpin this self-assessment.
1) Unconscious level: When a query is received, the brain initiates implicit memory retrieval, evaluating whether candidate memories fit the contextual cues. This process activates regions such as the orbitofrontal and prefrontal cortices within 300–500 milliseconds (Schnider, 2001; Irak et al., 2019), distinct from areas responsible for linguistic output, such as the posterior temporal lobe (i.e., Wernicke’s area) (Binder, 2015). Several factors—such as the amount, accessibility, and vividness of retrieved information—act as internal cues signaling potential knowledge (Koriat, 1993).
The dual-process theory (Kahneman, 2011) offers a broader framing: rapid, intuitive processes dominate simple recall tasks, while complex, deliberative processes support tasks such as reasoning or problem-solving.
Self-awareness in LLMs: A perspective on hallucination detection
1) Before-generation: Approaches that attempt to detect potential hallucinations *before* the model generates an answer mirror the unconscious processes of humans. These methods, including our own, rely on internal indicators such as activation patterns or uncertainty proxies to determine knowledge sufficiency prior to verbalization (Mallen et al., 2022). Benchmarks like Mintaka (Sen et al., 2022) and ParaRel (Elazar et al., 2021) primarily test immediate factual recall, aligning well with this layer of self-assessment.
2) After-generation: Other approaches evaluate the model’s response after generation, resembling conscious-level reasoning in humans. These include multi-pass generation, self-consistency checks, and the integration of external tools such as retrievers (béchard2024reducinghallucinationstructuredoutputs; Manakul et al., 2023; Chen et al., 2024). Benchmarks requiring structured reasoning, such as MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2022), and ELI5 (Fan et al., 2019), are more aligned with this stage, as they demand deliberative thought processes rather than pure retrieval.
Importantly, not all hallucinations can be attributed to failures in self-awareness. For instance, hallucinations in open-book QA tasks may arise from comprehension or inference errors rather than epistemic uncertainty. Benchmarks like SQuAD (Rajpurkar et al., 2016) and FEVER (Thorne et al., 2018) exemplify this issue, being more about information grounding than self-assessment.
@@ 3 Definition
In this section, we first examine the definition of self-awareness of human. Next, we define the task formulation of hallucination prediction, to establish a definition of self-awareness in LLMs. And we review the definitions from previous works. More detailed review on the related works of neuropsychology and LLMs is in §2.
Self-awareness of human
@@
In psychology, human self-awareness is defined as the capability of perception of one’s own mental processes or states, which includes thoughts, feelings, emotions, and knowing (Morin, 2011). In this work, we focus on the self-awareness of knowing certain knowledge, which is also referred to as self-knowledge (Yin et al., 2023).
@@
Some studies on LLMs also borrow the term “meta-cognition” from psychology to refer to self-knowledge (Li, 2023), which may be inaccurate. It is because meta-cognition focuses on the conscious level (Koriat, 1993), while recent research suggests that the human brain conducts judgment of knowing even at the unconscious level (Irak et al., 2019), which is even before consciously recognizing the meaning of the question.
Hallucination prediction
@@
The term “hallucination” has been broadly used to refer to the phenomenon where a model provides an incorrect answer to a given question (Li et al., 2023; Manakul et al., 2023; béchard2024reducinghallucinationstructuredoutputs). Thus, “hallucination detection” refers to the task of predicting whether a response is incorrect (Li et al., 2023; Chen et al., 2024).
@@
In this work, we focus specifically on (1) hallucination from factoid questions that examine whether the model possesses certain knowledge, as this has been widely used as a clear and straightforward scenario for exploring hallucination detection (Snyder et al., 2024; Zhang et al., 2024). (2) And we focus on an early detection (i.e., prediction) scenario (Snyder et al., 2024; Chen et al., 2024; Azaria and Mitchell, 2023) where $k$ is predicted before answer generation, as this setting is more appropriate for examine self-awareness (we describe this in the next section).
@@
To formulate the common problem setting of hallucination prediction, let $\theta$ represent the model, $x$ the query, and $y$ the answer label. The model $\theta$ infers $\hat{y}$ from the input $x$. The correctness of $\hat{y}$ can then be measured by evaluating its similarity to $y$, denoted as $k$, representing a binary value (True/False). Common evaluation methods include string matching (Zhang et al., 2024), GLUE, and G-eval (Liu et al., 2023). During this process, a datapoint $s_{\theta,x}=\theta(x)$ is extracted from $\theta$, which contains information about how $\theta$ perceived $x$. We denote this as $s$ for simplicity. Through a series of question-answering and evaluation processes, we obtain a dataset $\mathcal{D}=\{(s_{i},k_{i})\}_{i=1}^{N}$, where $N$ is the dataset size. From this, hallucination prediction is defined as a binary classification task where a learnable module $\phi$ learn to take input $s_{i}$ to predict $k_{i}$, which we note $\hat{k_{i}}=\phi(s_{i})$.
@@
Assuming $\theta$ as a transformer model (Vaswani et al., 2023), $s$ can be mainly two forms (1) Hidden state vectors: Transformer models are composed of multiple attention layers, where each layer passes a fixed-size vector (i.e., hidden states) to the next. These vectors encode the semantic and contextual information of the input $x$ (Reimers and Gurevych, 2019), and are also known to contain information about the model’s response that will be generated (Li et al., 2024a), providing a cue for hallucination prediction. (2) Confidence score: This refers to the softmax probability that the causal LLMs predict for the next token generation. While the hidden state is a high-dimensional representation (4096 dimensions in LLaMA-3-8B), the confidence score is a scalar value. As the hidden state contains richer information, it has achieved the best performance as a source for hallucination prediction and has been regarded as the main source (Snyder et al., 2024; Chen et al., 2024).
@@ 3.1 Formulating the self-awareness of LLM
We defined hallucination prediction as the process where $\phi$ perceives $s$ to infer $\hat{k}$ to predict $k$. As $s$ represents the state of the model after it has seen a question, it inherently contains information of two distinct objects, the question-side and the model-side, as illustrated in Figure 1. The question-side refers to the objective information that can be shared across different models, such as the domain of question (e.g., science, math, society) and the type of question (e.g., multiple choice, open-ended). For human, this type of information is derived from rational abilities (e.g., classification and reading comprehension), rather than from higher-order mental capacity such as self-awareness (Morin, 2011). In contrast, the model-side information refers to model-specific attributes, such as the possession of the knowledge that is needed for response, or the degree of confidence for answering. In humans, this corresponds to the domain of self-awareness.
Let us denote the representation of the question-side information in $s$ as $s_{Q}$, and the representation of the model-side information as $s_{M}$. We rewrite previously defined hallucination prediction $\hat{k}=\phi(s)$ as $\hat{k}=\phi(s_{Q},s_{M})$. When we denote the information contained in $s$ as $s$ itself, we can also note $s=s_{Q}\cup s_{M}$. This decomposition forms the basis for applying the Shapley value formulation.
Prior studies have also empirically shown that the hidden states of transformer models encode multiple properties in a linearly separable manner. For instance, Li et al. (2024a) demonstrates that the hidden state contains a direction associated with the attribute of “truthfulness”, therefore linearly adding this to the hidden states results in more truthful responses.
@@ 3.2 Definition from previous works
Utilizing $s$ as self-awareness Previous works implicitly regard self-awareness as the hallucination detection itself. In other words, the focus has been on the act of predicting $k$ from $s$, with no consideration given to the decomposition of $s$ into $s_{Q}$ and $s_{M}$. As a result, some of the hallucination detection performance reported in those works is partially overestimated by the effect of question-side shortcuts (Zhang et al., 2024; Azaria and Mitchell, 2023). We analyze such cases in §4.1.
Utilizing $s_{Q}$ as self-awareness In another case, some works define utilization of $s_{Q}$ as self-awareness, which is the opposite concept from our view. Yin et al. (2023) defines the term “self-knowledge” as a self-awareness on possession of certain knowledge. And from this definition, they construct the dataset SelfAware to measure the self-awareness capacity. While the definition in the paper is consistent with ours, the construction of the dataset stands on the opposite definition.
@@ 4.1 Case study on question-side shortcuts
We investigate sources of question-side shortcuts in datasets used in hallucination prediction studies. We focus on short-form closed-book factoid question answering scenarios, which include ParaRel (Elazar et al., 2021), Mintaka (Sen et al., 2022), HaluEval (Li et al., 2023), HotpotQA (Yang et al., 2018), and SimpleQuestion (Yin et al., 2016). We describe three main sources of question-side shortcuts, and further in §B.
(1) Broken Question Problem Many datasets have incomplete annotations for one-to-many question–answer pairs (e.g., Daniel Bernoulli → only “physics” labeled). This causes correct answers outside the label set to be marked as hallucinations, making the predictor $\phi$ learn domain-based shortcuts rather than true self-awareness. Datasets like Mintaka fix this by adding constraints for one-to-one mappings.
(2) Domain Shortcut The likelihood of hallucination (k=True) often varies by domain. For instance, if a model is strong in science but weak in history, history questions will naturally show higher hallucination rates. Thus, $\phi$ may simply learn domain-based correlations — predicting (k) based on what the question is about rather than what the model actually knows. This makes the task “question-aware” instead of “model-aware,” undermining the goal of self-awareness. A truly self-aware model should, like a human, recognize when it specifically knows or doesn’t know something, even within an unfamiliar domain.
(3) Question Type Shortcut The structure of the question itself (e.g., binary-choice, multiple-choice, open-ended) also influences (k). Binary-choice questions have a higher baseline chance of being correct $p(k=True)\geq 0.5$ even with random guessing, unlike open-ended ones $p(k=True)=0$. Consequently, $\phi$ might exploit this and always predict “True” for binary-choice questions, forming another non-self-aware shortcut. Datasets such as HotpotQA, HaluEval, and Mintaka contain such biases.